### v1

# Anomaly Induced Inflation in the Minimal Quartic Extension of Einstein’s Gravity

Abstract
A particular extension of Einstein’s General Relativity up to and including quartic terms in the curvature tensor is minimal in the sense that it has a unique maximally symmetric vacuum and only a massless spin-2 excitation in its spectrum around the vacuum. We study the inflation phase of the universe in the minimal quartic extension of Einstein’s gravity in the presence of trace anomaly terms in the Standard Model and the Minimal Supersymmetric Standard Model and show that the theory allows a quasi-de Sitter phase with a sufficient number of $e$-foldings and is compatible with the spectral index and the tensor to scalar ratio.
1 Introduction
The extreme homogeneity and the isotropy in the Cosmic Microwave Background (CMB) data [1] can be understood if the universe had gone through a brief period of accelerated expansion phase, or if the universe were a lot older. The latter option is unlikely. Within General Relativity such an expansion requires the existence of additional scalar fields, such as ”the inflaton” that first inflates the universe, gets affected by the inflation, and creates friction to gracefully exit the inflation phase. Another possible path to inflation is to modify the underlying gravity theory as was done by Starobinsky [2]. Here we follow this second route and study the inflationary phase of a modified gravity theory built on the curvature tensors up to and including the quartic terms. The theory was initially obtained as a subset of Born-Infeld-type gravity theory that was judiciously constructed to have all the good features of Einstein’s theory, such as the existence of a unique maximally symmetric vacuum and the non-existence of degrees of freedom besides a massless spin-2 graviton, but is much better behaved in the ultraviolet region of gravity. Here we first describe the theory before we study its cosmological solutions.
The Born-Infeld (BI) type gravity theories are a class of theoretical frameworks in gravitational physics that extend Einstein’s General Theory of Relativity (GR). They emerge from the need to address the inherent singularities present in GR, such as those associated with black holes and the Big Bang, by introducing higher derivative, and higher curvature corrections to the gravitational action. These corrections are inspired by the Born-Infeld electrodynamics (to be historically accurate these theories were first considered by Schrodinger and later by Eddington in gravity), originally formulated to regularize the self-energy of point charges. BI gravity theories have been studied in the context of string theory and modified gravity theories, offering new perspectives on the fundamental nature of gravity and its behavior in extreme conditions. The general form of the action of the BI type gravity theories can be given by
$I=\frac{1}{2\kappa_{0}\gamma}\int d^{4}x\sqrt{-\operatorname{det}\left(g_{\mu\nu}+\gamma A_{\mu\nu}\right)},$ (1)
where $\gamma$ is the Born-Infeld parameter and $A_{\mu\nu}$ is a second rank tensor constructed from the curvatures. Of course a proper $A_{\mu\nu}$ must be found that is consistent with all known tests of gravity and that also does not lead to inconsistencies, such as ghosts in the theory. For example, the naive, obvious choice $A_{\mu\nu}=R_{\mu\nu}$ leads to massive spin-2 ghosts in the spectrum. How ghosts can be avoided in these types of theories is a rather complicated problem, especially about a de-Sitter or anti-de-Sitter background. But fortunately, this problem has been fully solved which we discuss below.
Recently, a type of BI gravity model has been proposed [3, 4, 5, 7, 8] with a unique maximally symmetric solution and only a single massless spin-2 excitation about this vacuum, as a minimal extension of Einstein’s gravity. With a bare cosmological constant $\Lambda_{0}$, the action of the model is
$I=\frac{1}{2\kappa_{0}\gamma}\int d^{4}x\left(\sqrt{-\operatorname{det}\left(g_{\mu\nu}+4\gamma A_{\mu\nu}\right)}-\left(4\gamma\Lambda_{0}+1\right)\sqrt{-\operatorname{det}g}\right)$ (2)
where the second-rank tensor $A_{\mu\nu}$ has 3 undetermined constants, $a,b,c$ and is given as
$A_{\mu\nu}$ $=R_{\mu\nu}+cS_{\mu\nu}$
$+4\gamma\left(aC_{\mu\rho\nu\sigma}R^{\rho\sigma}+\frac{c+1}{4}R_{\mu\rho}R_{\nu}^{\rho}+\left(\frac{c(c+2)}{2}-2-b\right)S_{\mu\rho}S_{\nu}^{\rho}\right)$
$+\gamma g_{\mu\nu}\left(\frac{9}{8}C_{\rho\sigma\lambda\gamma}C^{\rho\sigma\lambda\gamma}-\frac{c}{4}R_{\rho\sigma}R^{\rho\sigma}+bS_{\rho\sigma}S^{\rho\sigma}\right).$ (3)
Here, $R_{\mu\nu}$ is the Ricci tensor while $S_{\mu\nu}$ is the traceless Ricci tensor, and $C_{\mu\rho\nu\sigma}$ is the Weyl tensor. The features, formulation of this theory, and its extension to generic $n$ dimensions were laid out in detail in [4], which relies on the foundation given in [5]. Note that this theory is in a sense the minimal BI gravity extension of Einstein’s theory as one can add more powers of curvature inside the determinant to obtain non-minimal extensions. Note also that as far as these theories are concerned, the $2+1$ dimensional BI gravity is rather unique, describing a massive graviton (instead of a massless one) with a remarkably simple Lagrangian of the form ${\mathcal{L}}=\sqrt{-\det\left(g_{\mu\nu}+\gamma G_{\mu\nu}\right)}$ [6] where $G_{\mu\nu}$ is the Einstein tensor. The success of this particular theory led us to search for similar theories in four and higher dimensions. Let us recapitulate some important aspects of the theory (2) that arise from the properties of the determinant and are not shared by most higher derivative gravity theories:
  • The theory has only a single massless spin-2 particle in the spectrum about its flat or (anti)-de Sitter vacuum.
  • The effective (dimensionless) cosmological constant ($\lambda$) of the maximally symmetric solution, i.e. $R^{\mu\nu}_{\,\sigma\rho}=\frac{\lambda}{3\gamma}(\delta^{\mu}_{\sigma}\delta^{\nu}_{\rho}-\delta^{\mu}_{\rho}\delta^{\nu}_{\sigma})$, obeys the quartic equation
    $4\lambda^{4}+4\lambda^{3}-\lambda+\lambda_{0}=0,$ (4)
    which has only one viable solution, i.e. the maximally symmetric vacuum is unique.
  • The effective Newton’s constant is related to the bare one and $\lambda$ via
    $\kappa=\frac{\kappa_{0}}{\left(1-4\lambda\right)\left(1+2\lambda\right)^{2}}.$ (5)
  • When expanded in small curvature ($|\gamma R|\ll 1$), the theory reproduces General Relativity as demanded by low energy observations at the first order; and at the quadratic order, it reproduces GR modified with the Gauss-Bonnet term, which at the classical level is just GR due to the topological nature of the Gauss-Bonnet term in four dimensions. The specific cubic and quartic order theories that we shall work with here were studied recently in [8] in the context of black hole physics.
2 Cosmological solutions
To investigate the cosmological solutions of the theory, let us consider the simple spatially flat Robertson-Walker metric,
$ds^{2}=-dt^{2}+a(t)^{2}d\vec{x}\cdot d\vec{x}.$ (6)
The field equation coming from the action is highly complicated as given in equation (F5) of [3]. Therefore we shall not directly use these field equations to search for solutions, but instead, we shall employ the symmetry reduction technique in the action before doing calculus of variations. The conventional approach involves inserting the metric into the action and defining the Hubble parameter as $H(t)=\frac{\dot{a}}{a}$ to simplify the Lagrangian into a single equation involving the Hubble parameter and its time derivatives. However, it is crucial to exercise caution in this reduction process, because the ”integrated-out” symmetry group in this case is not compact and, therefore, it falls outside the scope of Palais’ symmetric criticality guarantee [9, 10]. Consequently, merely varying the action concerning the scale factor or the Hubble parameter provides necessary but insufficient conditions. In this context, we adhere to Weinberg’s robust and error-free computation [11] for guidance. Inserting (6) into (2) and dropping an overall irrelevant constant, we arrive at the reduced action
[…]
where the scale factor $a$ and the Hubble parameter $H$ depend on the coordinate time $t$. The remarkable property of our theory is that just like Einstein’s gravity the second-time derivative of the Hubble parameter does not appear in the reduced action in contrast to other higher derivative theories. According to our action (2), the function ${\cal I}(H,\dot{H})$ can be written as
${\cal I}(H,\dot{H})=-1-4\lambda_{0}+\sqrt{X_{1}^{3}X_{2}},$ (8)
where $X_{1}$ and $X_{2}$ are given as
$X_{1}=2\dot{h}^{2}((c-2)c+4b-2)-2(c-2)\left(6h^{2}+1\right)\dot{h}+\left(6h^{2}+1\right)^{2}\,,$ (9)
$X_{2}=6\dot{h}^{2}(c(3c+10)-4b-6)+\left(6h^{2}+1\right)^{2}+6(c+2)\left(6h^{2}+1\right)\dot{h}\,,$
where we used the dimensionless Hubble parameter $h(t)\equiv\sqrt{\gamma}H(t)$ and the dimensionless time $\tilde{t}\equiv t/\sqrt{\gamma}$, so that derivatives are with respect to dimensionless time in (9).
It is clear that for the RW metric, the Weyl tensor vanishes identically and hence the parameter $a$ does not contribute to the reduced action. Note also that for the particular case $c=-1$ and $b=-5/2$, one has a purely quartic theory since $X_{1}=X_{2}$ of which the spherically symmetric black hole solutions were studied in [8]. Equipped with this, the necessary and sufficient equation to solve is the one coming from
${\frac{\delta I}{\delta g_{00}}\,\vline}_{RW}=0,$ (10)
which yields [11]
$\mathcal{I}-H\frac{\partial\mathcal{I}}{\partial H}+\left(-\dot{H}+3H^{2}\right)\frac{\partial\mathcal{I}}{\partial\dot{H}}+H\frac{d}{dt}\left(\frac{\partial\mathcal{I}}{\partial\dot{H}}\right)=0.$ (11)
By inserting the function $\cal{I}$ in (7) into the last equation, that is the EoM, which is the $00$-component of the corresponding field equations, can be expressed as follows.
$3H^{2}$ $=324\gamma^{3}\left(H^{8}+9H^{4}\dot{H}^{2}+6H^{3}\ddot{H}\dot{H}-H^{2}\dot{H}^{3}-3H\ddot{H}\dot{H}^{2}+\frac{3}{4}\dot{H}^{4}\right)$ (12)
$+108\gamma^{2}\left(H^{6}+\frac{9}{2}H^{2}\dot{H}^{2}+3H\ddot{H}\dot{H}-\dot{H}^{3}\right)+\Lambda_{0}\equiv\kappa_{0}\rho_{\text{MG}}\,,$
where we have kept the $00$-component of the Einstein tensor on the left, while we defined the rest coming from the modified gravity as $\rho_{\text{MG}}$ and wrote it on the right-hand side, as is often exercised in modified gravity theories. Moreover, from the variation of the action (2) with respect to the space component of the metric $\delta g_{ii}$ where $i=1,2,3$, or equivalently from the conversation law $\dot{\rho}_{\mathrm{MG}}+3H\left(\rho_{\mathrm{MG}}+p_{\mathrm{MG}}\right)=0$, one can find the pressure equation.
$-\left(2\dot{H}+3H^{2}\right)$ $=-324\gamma^{3}\left[H^{8}+\frac{8}{3}H^{6}\dot{H}+9H^{4}\dot{H}^{2}+11H^{2}\dot{H}^{3}+\frac{1}{12}\dot{H}^{4}\right.$ (13)
$\left.+2H\ddot{H}\dot{H}\left(6H^{2}+\dot{H}\right)+2\ddot{H}^{2}\left(H^{2}-\dot{H}\right)+\dot{H}\dddot{H}\left(2H^{2}-\dot{H}\right)\right]$
$-108\gamma^{2}\left[H^{6}+2H\dot{H}\left(H^{3}+\frac{9}{4}H\dot{H}+3\ddot{H}\right)+\ddot{H}^{2}+2\dot{H}^{3}+\dot{H}\dddot{H}\right]$
$-\Lambda_{0}\equiv\kappa_{0}p_{\text{MG}}\,,$
where we again used the same convention by defining $p_{\mathrm{MG}}$. Then, with the definition of $\rho_{\mathrm{MG}}$ and $p_{\mathrm{MG}}$, the field equations can be expressed in the usual form for the metric (6),
$\frac{3}{\kappa_{0}}H^{2}=\rho_{\text{MG}}\,,\qquad-\frac{1}{\kappa_{0}}\left(2\dot{H}+3H^{2}\right)=p_{\text{MG}}\,,$ (14)
which are called Friedmann equations. Notice that in the de Sitter limit, where the Hubble parameter is a constant $H(t)=H_{0}$, the equation of state (EoS) parameter tends to approach minus one,
${\omega_{0_{\text{MG}}}}\equiv\left.{\frac{p_{\text{MG}}}{\rho_{\text{MG}}}\,}\right|_{dS}=-1\,$ (15)
irrespective of the presence of a bare cosmological constant. The de Sitter solution can be found using the first Friedmann equation in (14).
$324H_{0}^{8}\gamma^{4}+108H_{0}^{6}\gamma^{3}-3H_{0}^{2}\gamma+\Lambda_{0}\gamma=0,$ (16)
which can be rewritten in dimensionless form as
$4\lambda^{4}+4\lambda^{3}-\lambda+\lambda_{0}=0,$ (17)
where we defined an effective cosmological constant $\Lambda=3H_{0}^{2}$ and corresponding dimensionless parameters $\lambda=\gamma\Lambda$ and $\lambda_{0}=\gamma\Lambda_{0}$ for effective and bare one respectively. Note that the last equation is (4). This polynomial equation has four solutions, two of which are imaginary and one of them is not physical in the sense that it gives $\lambda>1/4$ which corresponds to $\kappa<0$ according to (5) so that we have a unique de Sitter solution for given $\lambda_{0}$.
Here, we provide some remarks concerning the free parameters and the higher-order curvature terms within the Lagrangian. As mentioned previously, setting $a=0$, $b=-5/2$, and $c=-1$, the theory assumes a purely quartic nature, resulting in a Lagrangian with a finite number of curvature terms, specifically up to and including $R^{4}$ order terms. Alternatively, if we opt to keep these parameters as free variables, the conventional approach to derive field equations involves expanding $\operatorname{det}(\delta^{\nu}_{\mu}+4\gamma A_{\mu}^{\nu})$ and its square root in the small $\gamma$ limit. Consequently, the Lagrangian incorporates an infinite number of higher-order terms. However, in our analysis we found that the parameters $a$, $b$ and $c$ do not appear in the de Sitter equation and terms beyond the fourth order in curvature exhibit no influence on both the de Sitter solution and the linear perturbations around it. This observation implies that, when considering the simplified cosmological scenarios of a de Sitter universe and the quasi-de Sitter analysis, it is sufficient to focus on the lower-order curvature terms, thereby we can safely set the parameters to the desired values. In that case, the Lagrangian in (2) can be reduced to the simpler form.
$\mathcal{L}=\frac{1}{2\gamma\kappa_{0}}\left(\gamma\left(R-2\Lambda_{0}\right)+\gamma^{2}\frac{9}{2}\mathcal{G}+\gamma^{3}\frac{1}{2}\left(9\mathcal{G}R-R^{3}\right)+\gamma^{4}\frac{1}{8}\left(R^{4}-18\mathcal{G}R^{2}+81\mathcal{G}^{2}\right)\right)\,,$ (18)
where $\mathcal{G}\equiv R_{\mu\nu\sigma\rho}R^{\mu\nu\sigma\rho}-R_{\mu\nu}R^{\mu\nu}+R^{2}$ is the Gauss-Bonnet invariant. Then the BI gravity reduces to the form a $\mathcal{F}(R,\mathcal{G})$ theory with the action
$I=\frac{1}{2\kappa_{0}}\int d^{4}x\sqrt{-g}\mathcal{F}(R,\mathcal{G})\,,$ (19)
with a specific $\mathcal{F}$ given as
$2\gamma\mathcal{F}\equiv\left(1+\gamma R-\frac{1}{2}\gamma^{2}\left(R^{2}-9\mathcal{G}\right)\right)^{2}-4\gamma\Lambda_{0}-1\,.$ (20)
This is the minimal quartic extension of Einstein’s gravity in the sense that it has the same particle spectrum and a unique vacuum. With our parameter settings, the explicit form of Lagrangian (18) shows that it is quartic in curvature. Henceforth, in the remainder of the analysis, we shall adopt this simpler theory. In principle, one would like to study the process of inflation within the BI gravity described so far, but it turns out the theory does not have a quasi-de sitter phase. Such an attempt was made by one of us in [12] expecting that the theory would start from the unstable $\kappa<0$, vacuum, that is the repulsive gravity phase, and end up in the $\kappa>0$ vacuum in an inflated state. But, due to a computational error in that work, the claimed result is not correct. Hence, we must resort to the trace anomaly to realize inflation. We describe some basics of the trace anomaly in the next section before applying it to quartic theory.
3 Trace anomaly driven inflation in the minimal quartic extension of Einstein’s gravity
3.1 Trace anomaly
In one of the earliest models of inflation [2], the expansion is driven by the trace anomaly arising from a significant number of matter fields. The standard model of particle physics encompasses almost a hundred such fields, and this count at least doubles if we consider embedding the standard model within a supersymmetric framework. Consequently, a substantial number of matter fields existed in the early universe. At the energy scale of inflation, these fields can be considered as non-interacting and massless. Within this approach, we conduct a path integral over the matter fields, $\Phi$, within a specified background $g_{\mu\nu}$, yielding an effective action that is functional of the background metric:
$e^{i\Gamma\left(g_{\mu\nu}\right)}=\int\mathcal{D}\Phi e^{iS\left(\Phi,g_{\mu\nu}\right)}\,,$ (21)
where the classical action can be divided into vacuum and matter parts, $S(\Phi,g_{\mu\nu})=S_{\text{vac}}(g_{\mu\nu})+S_{\text{matter}}(\Phi,g_{\mu\nu})$. For a viable theory that is renormalizable, in addition to the Einstein-Hilbert action, the vacuum part should include four different four-derivative terms,
$S_{HD}=\int d^{4}x\sqrt{g}\left(b_{1}W+b_{2}\mathcal{G}+b_{3}\square R+b_{4}R^{2}\right)\,,$ (22)
which are called higher derivative terms that arise naturally in renormalization [13]. Then, the vacuum action is given by
$S_{\text{vac}}=S_{bac.}+S_{HD}\,.$ (23)
where $S_{bac}$ is the action of background gravitational model. Furthermore, the effective action of gravity $\Gamma[g_{\mu\nu}]$ admits a loop (or $\hbar$) expansion
$\Gamma\left[g_{\mu\nu}\right]=S_{vac}\left[g_{\mu\nu}\right]+\bar{\Gamma}^{(1)}+\bar{\Gamma}^{(2)}+\bar{\Gamma}^{(3)}+\ldots,$ (24)
so that one can write the total expression for the divergent part of the one-loop effective action of the vacuum for the theory involving $N_{S}$ real scalars, $N_{F}$ Dirac spinors and $N_{V}$ massless vectors
$\bar{\Gamma}^{(1)}_{div}=\frac{1}{\varepsilon}\int d^{4}x\sqrt{g}\left(\beta_{1}W-\beta_{2}\mathcal{G}+\beta_{3}\square R+\beta_{4}R^{2}\right),$ (25)
where $\beta$s are renormalization group $\beta$-functions that depend on the number of fields present in the theory and given as [14, 15]
$\beta_{1}=\frac{1}{120(4\pi)^{2}}\left(N_{\mathrm{S}}+6N_{\mathrm{F}}+12N_{\mathrm{V}}\right),\,\,\,\,\beta_{2}=\frac{1}{360(4\pi)^{2}}\left(N_{\mathrm{S}}+11N_{\mathrm{F}}+62N_{\mathrm{V}}\right),$ (26)
$\beta_{3}=\frac{1}{180(4\pi)^{2}}\left(N_{\mathrm{S}}+6N_{\mathrm{F}}-18N_{\mathrm{V}}\right),\,\,\,\,\beta_{4}=\frac{1}{2(4\pi)^{2}}N_{\mathrm{S}}\left(\xi-\frac{1}{6}\right)^{2}.$
Here, $\xi$ is the non-minimal coupling parameter for the term $\xi R\phi^{2}$ of a scalar field coupled to gravity, and the value $\xi=1/6$ corresponds to the special version of the scalar theory which possesses local conformal symmetry. Thus, in the conformal case, one can set $\beta_{4}=0$, then the action $S_{HD}$ satisfies the conformal Noether identity
$-\frac{2}{\sqrt{-g}}g_{\mu\nu}\frac{\delta S_{HD}}{\delta g_{\mu\nu}}=0\,,$ (27)
which means zero trace for the stress tensor of the vacuum $T^{\mu}_{\mu}=0$. At the quantum level, this condition is violated by an anomaly, the trace anomaly that reads as
$\left\langle T_{\mu}^{\mu}\right\rangle=-\frac{2}{\sqrt{-g}}g_{\mu\nu}\frac{\delta\bar{\Gamma}^{(1)}}{\delta g_{\mu\nu}}=\alpha W-\beta\mathcal{G}+\delta\square R\,,$ (28)
where $\alpha=\beta_{1}$, $\beta=\beta_{2}$, and $\delta=\beta_{3}$ for global conformal symmetry. Here, the emergence of the $\Box R$-term in the trace of the energy-momentum tensor is especially important since this is the term responsible for the Starobinsky instability. We should note that the coefficients $\beta_{1}$ and $\beta_{2}$ are independent of the renormalization scheme, but $\beta_{3}$ is not. For example, the notion adopted here that is the result given by the zeta-function regularization or predicted by AdS/CFT gives $-18$ as the coefficient of $N_{V}$ in the expression of $\beta_{3}$ while the result given by dimensional regularization has $+12$. The renormalization-dependent nature of $\delta$ suggests the possibility of adding a finite $R^{2}$ counterterm to the action at the classical level. When $\mathcal{N}=4$ super Yang-Mills theory is considered, for example, the field content can be given by a single parameter $N$ which is a large number. In that case, the effect of $\Box R$ vanishes and then the inflation never ends. However, $\delta$ can be adjusted to any desired value by adding the finite counterterm.
$S_{ct}=\frac{N^{2}\delta}{12(4\pi)^{2}}\int_{\mathcal{M}}d^{4}x\sqrt{-g}R^{2}\,.$ (29)
This counter-term explicitly breaks conformal invariance. In the literature, there is an ambiguity in the parameter $\delta$ and a comprehensive examination of this matter is available in [14, 16, 17]. The presence of $S_{ct}$ implies that we are effectively dealing with a higher derivative theory of gravity but it is arbitrary to consider it as part of the gravitational action or as part of the matter action. Since we are using the Born-Infeld action as the gravitational sector, the origin of the parameter $\delta$ can be regarded as a matter contribution. Hence, we provide particle content of three example models in Table 1, and adopted the Minimum Standard Model in our analysis mainly.
Model $N_{S}$, $N_{F}$, $N_{V}$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$
MSM (4, 24, 12) $73/(480\pi^{2})$ $253/(1440\pi^{2})$ $-17/(720\pi^{2})$
MSSM (104, 32, 12) $11/(48\pi^{2})$ $5/(24\pi^{2})$ $1/(36\pi^{2})$
[…]
Table 1: Particle contents of Minimum Standard Model (MSM), Minimal Supersymmetric Standard Model (MSSM) and $SU(\mathcal{N})$ Super Yang-Mills Theory, and corresponding values of $\beta$-functions.
Finally, the scenario can be succinctly encapsulated in the following manner: In curved space-time, the action governing the massless matter fields, that are scalars, Dirac spinors, and vectors, exhibits conformal invariance, yet the presence of one-loop vacuum contributions introduces certain divergences. Within the context of renormalization, specific counterterms are introduced to remove the singularities within the divergent part, albeit at the cost of breaking the conformal invariance inherent to the matter action itself. Classically, within a conformally invariant theory, the trace of the energy-momentum tensor is expected to be zero. However, the renormalization process results in the emergence of an anomalous trace of the energy-momentum tensor in (28), commonly referred to as the conformal anomaly or the quantum anomaly.
3.2 Trace anomaly and minimal quartic theory
The final action in the presence of quantum contributions (QC),
$I=\frac{1}{2\kappa_{0}}\int_{\mathcal{M}}d^{4}x\sqrt{-g}\left(\mathcal{F}(R,\mathcal{G})+2\kappa_{0}\mathcal{L}_{QC}\right)$ (30)
has the field equations:
$\mathcal{E}_{\mu\nu}\equiv$ $\mathcal{F}_{R}R_{\mu\nu}+\frac{1}{2}g_{\mu\nu}\left(\mathcal{G}\mathcal{F}_{\mathcal{G}}-\mathcal{F}\right)+\left(g_{\mu\nu}\square-\nabla_{\mu}\nabla_{\nu}\right)\mathcal{F}_{R}$ (31)
$+4\left[\left(2C_{\mu\sigma\nu\lambda}-R_{\mu\sigma\nu\lambda}\right)\nabla^{\sigma}\nabla^{\lambda}+\frac{R}{6}\left(g_{\mu\nu}\square-\nabla_{\mu}\nabla_{\nu}\right)\right]\mathcal{F}_{\mathcal{G}}=\kappa_{0}\left\langle T_{\mu\nu}\right\rangle$
where $\mathcal{F}$ is defined in (20), and $\mathcal{F}_{R}$ and $\mathcal{F}_{\mathcal{G}}$ are partial derivative of $\mathcal{F}(R,\mathcal{G})$ with respect to $R$ and $\mathcal{G}$. The energy-momentum tensor $\left\langle T_{\mu\nu}\right\rangle=\left\langle 0\left|T_{\mu\nu}\right|0\right\rangle$ emerges from the vacuum expectation value of the quantum corrections. Then, by taking trace of the field equations, we have
$R\mathcal{F}_{R}+2\mathcal{G}\mathcal{F}_{\mathcal{G}}-2\mathcal{F}+3\square\mathcal{F}_{R}-4G_{\mu\nu}\nabla^{\mu}\nabla^{\nu}\mathcal{F}_{\mathcal{G}}=\kappa_{0}\left\langle T_{\mu}^{\mu}\right\rangle\,.$ (32)
In the flat RW background with the metric (6), let us build the energy-momentum tensor as
$\left\langle T_{00}\right\rangle=\rho,\quad\left\langle T_{ij}\right\rangle=pa(t)^{2}\delta_{ij},\quad(i,j=1,2,3).$ (33)
Then, the field equations can be written in a simple form
$\begin{array}[]{r}\frac{3}{\kappa_{0}}H^{2}=\rho_{\mathrm{MG}}+\rho\equiv\rho_{\mathrm{eff}},\\ -\frac{1}{\kappa_{0}}\left(2\dot{H}+3H^{2}\right)=p_{\mathrm{MG}}+p\equiv p_{\mathrm{eff}},\end{array}$ (34)
where $\rho_{\mathrm{MG}}$ and $p_{\mathrm{MG}}$ are previously defined in Eqs. (12) and (13), and $\rho_{\mathrm{eff}}$ and $p_{\mathrm{eff}}$ are the effective energy density and pressure. In other words, these are the total energy constituents of the universe, comprising quantum matter components and modification of gravity. Furthermore, considering the covariant derivative of the field equations (31), $\nabla^{\mu}\mathcal{E}_{\mu 0}=0$, one arrives at the conservation law.
$\dot{\rho}+3H\left(\rho+p\right)=0\,.$ (35)
Now, we can obtain the full effective energy-momentum tensor. In principle, the complete structure of the energy-momentum tensor originating from the quantum corrections remains unknown. What we do possess knowledge of is its trace, denoted as $\left\langle T_{\mu}^{\mu}\right\rangle$. Nonetheless, when provided with the metric, it is possible to derive the energy-momentum tensor of the quantum anomaly on-shell. From the trace $\left\langle T_{\mu}^{\mu}\right\rangle$, we have
$-\rho+3p\equiv-\beta\mathcal{G}+\delta\square R,$ (36)
with $W=0$ in flat RW metric. Using the conservation law in Eq. (35) and eliminating the pressure, we can write
$\frac{d}{dt}\left(\rho a^{4}\right)=-\dot{a}a^{3}\left(-\rho+3p\right)=\equiv a^{3}\dot{a}\left(24\beta\frac{\dot{a}^{2}\ddot{a}}{a^{3}}+\delta(\ddot{R}+3H\dot{R})\right).$ (37)
The integration of this equation provides us the expression for the energy density
$\rho=\frac{\rho_{0}}{a^{4}}+6\beta H^{4}+\delta\left(18H^{2}\dot{H}+6\ddot{H}H-3\dot{H}^{2}\right),$ (38)
where $\rho_{0}$ is the integration constant. Using the conservation law, eventually one can find the pressure
$p=\frac{\rho_{0}}{3a^{4}}-\beta\left(6H^{4}+8H^{2}\dot{H}\right)-\delta\left(9\dot{H}^{2}+12H\ddot{H}+2\dddot{H}+18H^{2}\dot{H}\right).$ (39)
Note that the presence of $\rho_{0}$ implies that the quantum state may contain an arbitrary amount of radiation. However, we can reasonably disregard this radiation due to our focus on extremely high energy scales, all the way to the Planck scale. At this scale, the energy density associated with the standard radiation is typically insignificant compared to that of quantum corrections and gravitational modifications. Consequently, we set
$\rho_{0}=0\,.$
To study the cosmological evolution of the model, it suffices to use one of the field equations (34). In what follows, we use the first one, where $\rho_{\mathrm{MG}}$ is defined in (12) and $\rho$ is found in (38) with $\rho_{0}=0$.
3.3 de Sitter solutions
Let us recast the EoM which is the first equation in the Friedmann equations (34) in full form.
$3H^{2}=$ $\Lambda_{0}+3\kappa_{0}\left(2H^{4}\beta+\delta\left(6H^{2}\dot{H}+2H\ddot{H}-\dot{H}^{2}\right)\right)$ (40)
$+\gamma^{3}\left(324H^{8}+2916H^{4}\dot{H}^{2}+1944H^{3}\ddot{H}\dot{H}-324H^{2}\dot{H}^{3}-972H\ddot{H}\dot{H}^{2}+243\dot{H}^{4}\right)$
$+\gamma^{2}\left(108H^{6}+486H^{2}\dot{H}^{2}+324H\ddot{H}\dot{H}-108\dot{H}^{3}\right).$
Here, the terms with $\kappa_{0}$ are the effect of trace-anomaly while the effect of modified gravity appears as the terms multiplied by $\gamma$. Solution for a constant $H(t)=H_{0}$ gives us de Sitter solution which is the solution of the resultant polynomial equation,
$0\displaystyle 324H_{0}^{8}\gamma^{3}+108H_{0}^{6}\gamma^{2}+6H_{0}^{4}\kappa_{0}\beta-3H_{0}^{2}+\Lambda_{0}=$ 0 (41)
$\frac{2\beta_{0}\lambda^{2}}{3}+4\lambda^{4}+4\lambda^{3}-\lambda+\lambda_{0}=0$
where we defined $\beta_{0}=\frac{\beta\kappa_{0}}{\gamma}$ and $\delta_{0}=\frac{\delta\kappa_{0}}{\gamma^{2}}$ to be used later. The second line of (41) shows that only the Gauss-Bonnet term in the anomalous trace contributes to the de Sitter solution, on the other hand, we will see that the coefficient $\delta$ in front of $\square R$ term plays a crucial role in stability analysis of de Sitter solution. Here, the appearance of a term of $\lambda^{2}$ order makes a double de Sitter solution possible in the physical range $\kappa>0$, depending on the values of the parameters. One of them is due to the presence of the bare cosmological constant $\Lambda_{0}$, and the other is due to the trace anomaly, and if the latter is unstable it corresponds to inflation.
3.4 Dynamics of inflation
We first analyze our main differential equation (40) with a phase portrait analysis. Using a change of variables,
$x=H(t)\quad\text{and}\quad y=\dot{H}(t),$ (42)
it can be written as a first order differential equation
$\frac{dy}{dx}=f(x,y\,;\,\Lambda_{0},\gamma,\beta,\delta)=\frac{f_{1}(x,y)}{f_{2}(x,y)}\,.$ (43)
Here, $\Lambda_{0}$ and $\gamma$ are free parameters that need to be determined while the other parameters $\beta$ and $\delta$ are determined by the number of particles in the underlying particle theory. The functions $f_{1}(x,y)$ and $f_{2}(x,y)$ are
$f_{1}(x,y)$ $=\Lambda_{0}+324\gamma^{3}x^{8}+108\gamma^{2}x^{6}+6x^{4}\left(\beta\kappa_{0}+486\gamma^{3}y^{2}\right)$ (44)
$+x^{2}\left(-324\gamma^{3}y^{3}+486\gamma^{2}y^{2}+18\delta\kappa_{0}y-3\right)+243\gamma^{3}y^{4}-108\gamma^{2}y^{3}-3\delta\kappa_{0}y^{2}$
$f_{2}(x,y)$ $=6xy\left(\delta\kappa_{0}+54\gamma^{2}y\left(6\gamma x^{2}+1\right)-162\gamma^{3}y^{2}\right)$
With this reparametrization, the phase portrait of the system can be analyzed in the $xy$-plane as in Fig. 1. We provided the phase portrait for the Minimum Standard Model (MSM) in four different regimes.
In the first scenario depicted in Fig. 1 (a), we examine the case where there is no contribution from the modified gravity, representing a pure trace-anomaly situation within the framework of Einstein’s gravity. Two critical points can be seen along the $y=0$ axis, corresponding to two distinct de Sitter solutions. For the particle content of MSM, the second solution is identified as unstable. Consequently, the universe is presumed to have originated nearly from this point, having a negative $\dot{H}$, subsequently transitioning towards the first critical point, which acts as the final attractor of the system, determined by $\Lambda_{0}$, and may correspond to the late-time acceleration. In the absence of an intrinsic cosmological constant, the ultimate attractor point resides at $(x,y)=(0,0)$. Thus, during the inflationary epoch, the Hubble parameter $H$ experiences a slow decrease while maintaining $\dot{H}$ in near constancy, eventually oscillating around the final attractor, indicative of the reheating era.
(a) Einstein’s gravity with trace-anomaly. ($\gamma=0$)
(b) Minimal quartic model without trace-anomaly. ($\gamma=0.0025$)
(c) Minimal quartic model with trace-anomaly. ($\gamma=0.0025$)
(d) Minimal quartic model with trace-anomaly. ($\gamma=0.0060$)
Figure 1: The phase-portrait of the system described by the EoM 40 for the MSM case with parameter settings $x=H(t)$ and $y=\dot{H}(t)$. The figure is plotted in $c=\hbar=\kappa_{0}=1$ units and the value of the bare cosmological constant is chosen arbitrarily as $\Lambda_{0}=6$.
In the second scenario, depicted in Fig. 1 (b), we switch off the trace anomaly and investigate the dynamics of pure modified gravity. Once again, two distinct de Sitter solutions are evident along the $y=0$ axis. Additionally, a solid black line is introduced along the x-axis to indicate a non-physical region where $\lambda>1/4$, a characteristic feature typical of Born-Infeld-type theories, setting an upper limit on the effective cosmological constant. Consequently, the second solution is nonphysical, as it intersects with the solid line, resulting in the existence of a unique de Sitter solution. In this context, a critical surface manifests along the $y=0$ axis, changing the characteristics of critical points both below and above this surface. The unique de Sitter solution remains stable within the domain where $\dot{H}<0$, yet transitions to instability when $\dot{H}>0$ is encountered. Furthermore, the critical surface is identified as unstable within these points, while exhibiting attractor behavior beyond this region.
In Fig. 1 (c), we explore our modified gravity theory gently introducing a small value for $\gamma$ in the presence of the trace anomaly. In comparison to the first scenario, the introduction of modified gravity results in the emergence of three critical points situated above the de Sitter solutions observed in the original trace-anomaly case. However, one of these critical points exhibits a notably high $\dot{H}$ value, surpassing the upper limit of this plot and consequently playing no role in the inflationary solutions. The remaining two critical points reside above the critical points along the $y=0$ axis and are positioned on a critical surface represented by a curve that converges toward the x-axis as $H\rightarrow\infty$. Notably, the critical surface with critical points on it in pure modified gravity case is raised within the $\dot{H}>0$ region, while still maintaining the de Sitter solution due to the trace anomaly within the physically plausible range, $\kappa>0$. Under an initial condition where $\dot{H}\lesssim 0$, the dynamics of inflation remain largely unaffected by the presence of modified gravity.
In Fig. 1 (d), to understand the effect of modified gravity on trace-anomaly-driven inflation, we play with the value of $\gamma$ and increase the intensity it. Increasing it makes the above critical points closer to the de Sitter points, while the solid line indicating $\lambda>0$ region, started to cover up the critical point corresponding to inflation so that inflation cannot happen. This imposes an upper bound on $\gamma$ so that the effects of modified gravity should not exceed the trace anomaly for a viable inflationary model.
Furthermore, we investigate the scenario wherein the sign of the $\Box R$-term is positive, indicative of an initial stable point. This condition finds representation in the context of MSSM as an illustrative example model encompassing the particle content pertinent to this case. In Fig.2, the phase-portrait of the system is plotted without and with the effects of modified gravity for the MSSM case respectively, and from Fig.(a), it can be seen that the second de Sitter solution due to the trace-anomaly is a stable critical point. In this case, a graceful exit from Inflation requires its mechanism to disrupt this state. In Ref. [18], a possible transition from stable to unstable phases of inflation is discussed, and different quantum mechanisms that may produce large $R^{2}$-term during the inflation are proposed for a graceful exit for example. On the contrary, we show that the minimal extension of Einstein’s gravity makes possible exit from inflation without any other mechanism. In Fig.2(b), we have plotted the contribution of modified gravity in this case. As in MSM case, modified gravity creates additional critical points on the phase-portrait, visible ones of those in the plotted region are labeled as $P_{i}(x_{i},y_{i})$ in the graph, but this time the two of them lying on a critical surface are in $\dot{H}<0$ region. Note that increasing the intensity of modified gravity makes this critical surface closer to the $y=0$ line. Therefore, if the universe starts with $H\simeq x_{1}\simeq x_{2}$ and a small value of $\dot{H}$ from below this surface, $\dot{H}\lesssim y_{2}$ for example, then the exit from inflation can occur due to the attractor behavior of the surface between 00 and $x_{3}$. This prevents $\dot{H}$ from flowing through negative infinity and ends inflation with $H\lesssim x_{3}$. If the strength of modified gravity is high enough, the initial point for $\dot{H}$ can be in the slow-roll regime, i.e. $\left|\dot{H}/H^{2}\right|\ll 1$ while keeping $\kappa$ positive.
(a) Trace-anomaly in Einstein’s gravity with MSSM particle content
(b) Trace-anomaly in minimal quartic model with MSSM particle content
Figure 2: The phase-portrait of the system described by the EoM 40 for MSSM case. The figure is plotted in $c=\hbar=\kappa_{0}=1$ units and the parameters are set as $\gamma=0.004$ and $\Lambda_{0}=6$.
3.4.1 Perturbations around the de Sitter solution
To find the stability condition of the de Sitter solution, we analyze the linearized perturbations around it $H(t)=H_{0}+\Delta H(t)$. Substituting this into (40), and linearizing it in small $\Delta H(t)/H_{0}\ll 1$ gives a solution of the perturbation in the following form,
$\Delta H(t)=A_{0}\,e^{\xi H_{0}t},$ (45)
where $A_{0}\neq 0$ is the amplitude of perturbation; and $\xi$ is the solution of the quadratic equation,
$6H_{0}\left(-432H_{0}^{6}\gamma^{3}-108H_{0}^{4}\gamma^{2}-4H_{0}^{2}\gamma\beta_{0}-H_{0}^{2}\kappa_{0}\delta\xi^{2}-3H_{0}^{2}\kappa_{0}\delta\xi+1\right)=0,$ (46)
which can be expressed in different forms,
$\xi$ $=-\frac{3}{2}\pm\frac{\sqrt{\kappa_{0}\delta\left(-1728H_{0}^{6}\gamma^{3}-432H_{0}^{4}\gamma^{2}-16H_{0}^{2}\gamma\beta_{0}+9H_{0}^{2}\kappa_{0}\delta+4\right)}}{2H_{0}\kappa_{0}\delta}$
$=-\frac{3}{2}\pm\frac{\sqrt{\delta_{0}\left(-16\beta_{0}\lambda+9\delta_{0}\lambda-192\lambda^{3}-144\lambda^{2}+12\right)}}{2\delta_{0}\sqrt{\lambda}}$
$=-\frac{3}{2}\left[1\pm\sqrt{1+\frac{4}{3\delta}\left(\frac{1}{\kappa\Lambda}-\frac{4\beta}{3}\right)}\,\right].$ (47)
The last expression looks exactly the same as the expression with pure trace-anomaly driven inflation [19], but the difference is that $\Lambda$ is now the effective cosmological constant which is the solution to (41) with $\lambda=\gamma\Lambda$ and $\lambda_{0}=\gamma\Lambda_{0}$, and $\kappa$ is the effective gravitational constant whose relation with the bare $\kappa_{0}$ and $\Lambda$ is given in (5); so that $\kappa$ implicitly involves $\Lambda$ also. Furthermore, if inflation is described by an unstable de Sitter solution which requires $\xi>0$, the condition for viable inflation can be given as
$\kappa\Lambda>0\,\land\,\beta>\frac{3}{4\kappa\Lambda}\,\land\,\delta<0.$ (48)
Hence, the instability depends on the sign of the $\square R$ term in the trace of energy-momentum tensor. For the particle content of MSM, $\delta$ is negative, while MSSM gives positive values. As discussed above, the renormalization scheme dependence of this term makes it possible to adjust it to any value by introducing a counter $R^{2}$-term in Lagrangian. In most of the analysis, however, we simply consider the MSM case.
Inflation ends when the perturbation $H_{0}\left|\Delta H\right|$ is of the order of $H_{0}^{2}$, namely, with $A_{0}$ being a constant, $A_{0}\exp{\left(\xi H_{0}t_{f}\right)}\simeq H_{0}$, then the time when the inflation ends, $t_{f}$, can be estimated as
$t_{f}\simeq\frac{\zeta}{\xi H_{0}},$ (49)
where we have defined a new parameter $\zeta$ as
$\zeta=\log{\left(\frac{H_{0}}{\left|A_{0}\right|}\right)},$ (50)
since it will appear in the results frequently. If inflation is governed by a (quasi) de Sitter solution where $a(t)\simeq\exp{\left(H_{0}t\right)}$, we can introduce the number of e-folds $\mathcal{N}$ as
$\mathcal{N}\equiv\log{\left(\frac{a_{f}}{a_{i}}\right)}=H_{0}t_{f}.$ (51)
Moreover, we can define the effective EoS parameter for the effective fluid whose energy density and pressure is defined in (34).
$\omega_{\text{eff}}=\frac{p_{\text{eff}}}{\rho_{\text{eff}}}.$ (52)
For the de Sitter solution, $\omega_{\text{eff}}=-1$, but by introducing perturbations $H(t)=H_{0}+\Delta H(t)$, we obtain
$\omega_{\text{eff}}=-1-\frac{2\Delta\dot{H}}{3H^{2}}\simeq-1-\frac{2\xi A_{0}e^{\xi H_{0}t}}{3H_{0}}+\mathcal{O}(A_{0}^{2}).$ (53)
The accelerated expansion requires $w_{\text{eff}}<-1/3$. If the acceleration ends at $t_{f}$ in (49), by using the solution (45) and the number of $e$-folding $\mathcal{N}$, it can be said that during the inflation, the value of effective EoS changes between
$-1<\omega_{\text{eff}}\leq-1+\frac{2\zeta}{3\mathcal{N}}\,.$ (54)
Hence, to stop the acceleration at the end, i.e. $\omega_{\mathbf{eff}}(t_{f})\geq-1/3$, we found that the sign of $A_{0}$ is negative and the value should be $|A_{0}|\ll H_{0}$. Hence, the Universe starts from nearly de Sitter stage.
3.5 Observable constraints
Let us study some of the observables in the quartic theory. From the second time derivative of $a(t)$,
$\frac{\ddot{a}}{a}=H^{2}+\dot{H}=H^{2}(1-\epsilon),$ (55)
we can introduce the slow-roll parameter $\epsilon$, which has to be small during the inflation to vary $H(t)$ slowly.
$\epsilon=-\frac{\dot{H}}{H^{2}}\ll 1.$ (56)
To have viable inflation with a graceful exit, the slow-roll requires that the change in $\epsilon$ should also be small, so that we have another parameter $\eta$ to measure it.
$|\eta|=\left|-\frac{\ddot{H}}{2H\dot{H}}\right|\equiv\left|\epsilon-\frac{1}{2\epsilon H}\dot{\epsilon}\right|\ll 1.$ (57)
These two parameters, referred to as slow-roll parameters, play a pivotal role in analyzing the feasibility of inflation within the framework of standard inflationary models governed by scalar fields. In terms of these parameters, the current cosmological data constrain the spectral index $n_{s}$ and the tensor-to-scalar ratio $r$, which are given as
$n_{\mathrm{s}}=1-6\epsilon+2\eta\quad\text{and}\quad r=16\epsilon.$ (58)
Hence, in our model, we have the slow-roll parameters
$\epsilon\simeq-\frac{\Delta\dot{H}(t)}{H_{0}^{2}}=-\xi\frac{A_{0}}{H_{0}}e^{\xi H_{0}t}\,,$ (59)
$\eta\simeq\epsilon-\frac{\dot{\epsilon}}{2\epsilon H_{0}}=-\frac{\xi}{2}\left(1-\frac{A_{0}}{H_{0}}e^{\xi H_{0}t}\right)\,,$ (60)
and the spectral index
$n_{s}=1+\xi\left(7\frac{A_{0}}{H_{0}}e^{\xi H_{0}t}-1\right)+\mathcal{O}(A_{0}^{2})$ (61)
in terms of the model parameters. Since $|A_{0}|\ll H_{0}$, and $t\ll t_{f}$ during the inflation, we can express the spectral index approximately in a familiar form in terms of the number of $e$-folding,
$n_{s}\simeq 1-\frac{\zeta}{N}\,.$ (62)
Accordingly, the tensor-to-scalar ratio is given by
$r\simeq\frac{16\zeta e^{-\zeta}}{N}\,.$ (63)
To check if the model is consistent with the observations, we provide the observational constraints from the PLANCK satellite. According to the latest results [20] published in 2020,
$n_{s}=0.9649\pm 0.0042\;(68\%\text{CL})\quad\text{and}\quad r<0.101\;(95\%\text{CL})\,.$ (64)
The universe requires at least $N=60$ $e$-foldings during the inflation phase to address chronic problems such as flatness and horizon problems. Therefore, using 62 and according to Planck satellite results, we find that $\zeta\approx 2.4$. This yields the initial perturbation $A_{0}=H_{0}e^{-2.4}$, which is as small as required, and $\xi=0.04$ can be determined. Then, utilizing the second line in 47, we can establish the relationship between $\lambda$ and $\gamma$ within the allowed region of observations. In Figure 3, for instance, we depict the $\lambda$ vs $\gamma$ relationship and observe that for a small $\gamma<1$, there exists a value of $\lambda<1/4$ such that it falls within the physical range ($\kappa>0$) and is consistent with Planck results.
Figure 3: The relation between the effective dimensionless cosmological constant $\lambda$ vs the Born-Infeld parameter $\gamma$ that is consistent with Planck results ($n_{s}=0.96$) and in $c=\hbar=\kappa_{0}=1$ units.
In addition, with these parameter settings for example, the tensor-to-scalar ratio is found as $r\approx 0.058$ from the equation 63, which also satisfies observational constraints.
4 Conclusions
We have studied the cosmological inflation era in a theory that extends General Relativity in such a way that the resultant theory shares some nice properties of the former, such as the uniqueness of the maximally symmetric vacuum and the non-existence of degrees of freedom other than a massless spin-2 graviton while being a better-behaved theory in the ultraviolet regime of gravity. In principle, the theory is of the Born-Infeld type and includes infinitely many powers of curvature in small curvature expansion; but we studied a particular limit of the theory which has judiciously chosen curvature terms up to and including quartic order. The resulting theory becomes a particular $\mathcal{F}(R,\mathcal{G})$ theory where $R$ is the scalar curvature and $\mathcal{G}$ is the Gauss-Bonnet invariant. This theory when augmented with the inevitable trace anomaly as computed in the Standard Model and the Minimal Supersymmetric Standard Model allows a quasi-de Sitter solution with enough number of $e$-foldings and is consistent with the observed spectral index and the tensor to scalar ratio. In the Supersymmetric Model, the cubic and quartic terms allow an exit from the otherwise stable de Sitter phase if the initial value of the time derivative of the Hubble parameter is large enough while still being in the slow-roll regime.
5 Acknowledgment
We would like to thank Prof. Dr. Sakir Erkoc for their kind support.